OpenStudy (anonymous):

\sqrt{20 + x} is continuous on the interval: [-20, infinity), is this correct?

OpenStudy (turingtest):

yes

OpenStudy (alfie):

What is the domain of the square root? >= 0 So if you put -20 inside the square root what do you get? 0 Which is a valid point of the domain, so the function will be continuous in this interval.

OpenStudy (turingtest):

\[20+x \ge 0\rightarrow x \ge -20\]in interval notation\[[-20,\infty)\]

OpenStudy (anonymous):

yup